A) At what angle will 642 nm light produce a first-order maximum when falling on a grating whose slits are 1.8 × 10-3 cm apart?

B)At what angle will 513 nm light produce a second-order maximum when falling on a grating whose slits are 1.19 × 10-3 cm apart?

C) A grating that has 3,904 slits per cm produces a third-order fringe at a 21.7° angle. What wavelength of light is being used in nm?
Answer #1

Answers are: (a) 2004® (b) 4.95° (C) 316 nmQF (a) d = 642 nm z 642 x 10-9m d = in8x10-3 cm = 1.8x10-5 m m=1 for first order. dsino = md Pro- in da sin (md) o= sin ( 1xo= 4.95° 0=21070, N = 3904 slits/cm, n=3 d = 1 n 1304 = 2.56810 cm. dsino and de deine - (256x104) sin (21.7) W da 3016810-5

Hope this helps...